Science Report May 2nd

Hello mission control,

Here is the Science Report for 2 May 2018

Crew 193 – PHEnOM Gold Crew

Submitted by XO/Science Officer Doug Campbell

1. Emergency EVAC EVA: No EVAs occurred today due to weather.

2. Sunspot and Solar Flare Monitoring: a very small prominence was found on the sun today.

3. Shortwave Texting on EVA’s: No EVAs occurred today due to weather

4. GPS Route Measurements: No EVAs occurred today due to weather

5. Waterless dish cleaning: Bacteria growth and monitoring will continue over the remaining days of the mission. The study is starting to show promising results regarding the use of UV and a cloth to clean dishes on mars without the use of water. Using dirt to clean the dishes seems to have some rapid growing bacteria!

6. The MAG (Maximum Absorbency Garment) utilization study: The study has been scrapped. Results will be from blue crew only.

7. Spacesuit Helmet Fogging: Study has been discontinued after 7 days of data collection.

8. Crew Comfort, Health and Safety on EVA: No EVAs occurred today due to weather




Crew 193 XO and Science Officer

Copyright © The Mars Society. All rights reserved. | Main Site
%d bloggers like this: